A body projected vertically up crosses point A and B separated by 28 m with velocities one third and one fourth of the initial velocity; respectively.what is the maximum height reached by it above the ground.
1
Expert's answer
2020-03-27T10:46:58-0400

Let X represent initial speed. Then speed at point A is X/3 and speed at B is X/4.

Distance between A and B is equal to:

"Dab = ((X\/3)^2-(X\/4)^2)\/(2*g) = 28"

Solving the equation for X, obtaining X = 106.3 m/s.

Maximum height is determined as follows.

"Hmax=X^2\/(2*g) = 106.3^2\/(2*9.81) = 576 m"